### v1

@@ Abstract
WebAssembly (Wasm) programs may trigger bugs in their engine implementations. To aid debugging, program reduction techniques try to produce a smaller variant of the input program that still triggers the bug. However, existing execution-unaware program reduction techniques struggle with large and complex Wasm programs, because they rely on static information and apply syntactic transformations, while ignoring the valuable information offered by the input program’s execution behavior.
We present RR-Reduce and Hybrid-Reduce, novel execution-aware program reduction techniques that leverage execution behaviors via record and replay. RR-Reduce identifies a bug-triggering function as the target function, isolates that function from the rest of the program, and generates a reduced program that replays only the interactions between the target function and the rest of the program. Hybrid-Reduce combines a complementary execution-unaware reduction technique with RR-Reduce to further reduce program size.
We evaluate RR-Reduce and Hybrid-Reduce on 28 Wasm programs that trigger a diverse set of bugs in three engines. On average, RR-Reduce reduces the programs to 1.20% of their original size in 14.5 minutes, which outperforms the state of the art by 33.15 in terms of reduction time. Hybrid-Reduce reduces the programs to 0.13% of their original size in 3.5 hours, which outperforms the state of the art by 3.42 in terms of reduced program size and 2.26 in terms of reduction time. We envision RR-Reduce as the go-to tool for rapid, on-demand debugging in minutes, and Hybrid-Reduce for scenarios where developers require the smallest possible programs.
@@ I Introduction
We release our tools and data as open source, for others to reproduce and build on our results [18].
@@ III-A Problem Statement
To the best of our knowledge, we are the first to formulate and address the execution-aware program reduction problem. The key difference between the execution-aware program reduction problem and the standard program reduction problem is that the former considers an execution of the input program, while the latter does not.
@@ III-B Overview
Figure 3 shows an overview of RR-Reduce. Given an executable program as an input, RR-Reduce first heuristically identifies functions in the program that may be critical for triggering the bug, which the approach considers as candidates for the target function (Section III-D). Next, the approach splits the input program into two parts: one containing the target function and another containing all remaining functions (Section III-E). The resulting partitioned program is then passed to an existing record and replay technique, which records the execution and produces replay functions (Section III-F). Importantly, we configure the record and replay technique to focus the replay on replaying only the target function. That is, the remaining functions are either removed or replaced with replay functions that mimic those parts of the behavior of the original functions that are necessary to accurately replay the target function. Given the replay functions, RR-Reduce combines the target function with the replay functions to produce a candidate for the reduced program (Section III-G). Finally, the oracle, also given as an input, checks whether the candidate program still triggers the bug (Section III-H). If so, and if the candidate program is smaller than the input program, RR-Reduce returns the candidate program as the reduced program. Otherwise, the approach repeats the process with a different target function.
@@ III-C Running Example
Figure 4 shows a running example of RR-Reduce. On the left, we have an input program that consists of three functions: a, b, and c. Suppose an execution of this input program triggers a wrong-code bug in a Wasm engine, which results in a stack trace that contains the c function. Because c is in the stack trace, RR-Reduce heuristically identifies c as the target function. Next, RR-Reduce splits the input program into two parts: one containing the target function c and another containing the remaining functions a and b. Then, RR-Reduce uses an existing record and replay technique, Wasm-R3 (Section II-B), to produce replay functions that accurately replay the execution of the target function c. This is achieved in three phases: record, reduce, and replay. The resulting replay functions mimic the behavior of the remaining functions a and b in the input program. Our approach combines these replay functions with the target function c to produce a candidate program, and checks using the oracle whether the candidate program still triggers the bug in the Wasm engine and is smaller than the input program.
@@ IV-A Experimental Setup
We measure the size of a Wasm program in terms of its code size, which is the size of the code section of the Wasm program. This excludes the size of the data section, which contains initialization data for the memory, and the size of the custom section, which contains debug symbols. Focusing on the code size is motivated by the fact that the code section is what engine developers typically focus on when debugging engine bugs. We report two kinds of code sizes: the “All” code size, which includes all functions in the binary, and the “Target” code size, which is the size of the target function. The “Target” is most relevant for engine developers, as they typically need to understand the function that triggers the bug.
@@
We implement RR-Reduce through a combination of Python, Rust, and JavaScript, building on several libraries and tools in the Wasm ecosystem. The first step is to identify candidates for the target function (Section III-D). We use the existing tool, wasm-tools objdump [28], to compute the All set, a dynamic analysis based on the Wizard engine [13] to compute the Dynamic set, and implement the computation of the Heuristic set in Python. Next, our implementation splits the input program into two Wasm binaries (Section III-E), which we implement in Rust. In addition to two newly created Wasm binaries, our implementation also creates JavaScript glue code to dynamically link the two Wasm binaries. To resolve the circular dependency between the two binaries, the JavaScript code creates a closure that calls the target function, and instantiates the replay binary with this closure as an import. The record and replay step uses the existing Wasm-R3 [12] (commit hash e3fa5d4). Finally, our implementation combines the target function with the replay functions (Section III-G) using the wasm-merge tool [29]. To speed up the reduction process, our implementation tries to reduce the input program for different target functions in parallel. Our experiments are conducted on a machine running Ubuntu 24.04 LTS, equipped with an Intel Core i9-13900k CPU with 32 logical cores and 192GB of DRAM.
@@
We evaluate the effectiveness of our approach by applying RR-Reduce and Hybrid-Reduce, as well as the two baselines, to the Wasm programs in Table I. We run each technique either until it terminates or a timeout of 24 hours is reached. We run eight reduction tasks in parallel, allowing each task to use a maximum of four logical cores.
@@
The real strength of RR-Reduce is revealed when it is combined with existing program reduction techniques into a hybrid approach. Hybrid-Reduce reduces the input program to 0.13% of its original size, which is a 3.42% improvement over the current state of the art, wasm-reduce. Hybrid-Reduce achieves this by giving the best result in 23 cases, of which six are ties with the current state of the art. The benefit of the Hybrid-Reduce is not only in its better average performance, but also in its effectiveness on programs where the current state of the art struggles. Table III highlights the difference. Specifically, wasm-shrink produces programs larger than 100KB in 15 cases, and wasm-reduce does so in three cases. In contrast, all Hybrid-Reduce-reduced programs are well below 100KB, and often even further.
IV-C RQ2. Efficiency
IV-D RQ3. Qualitative Analysis
@@
Boa, which is 5.2MB in size, triggers a wrong-code bug111Fixed by https://github.com/titzer/wizard-engine/commit/6d2b057 that gets triggered when the interpreter tries to interpret a jump with an offset bigger than 32,768 bytes. Wasm-reduce reduces the input program to 317KB, which is still too large to manually debug. It also contains 87 functions besides the function where the jump happens, which further complicates debugging. Wasm-reduce cannot simply remove these functions because other functions are necessary for the jump to happen. In contrast, RR-Reduce takes a different approach: It deletes the bodies of 453 executed functions and replaces the bodies of eight functions with replay functions. Among the nine functions that are executed in the reduced program, seven are short functions that update the state, and one is the entry to the reduced program, which calls the bug-triggering functions. The remaining one is the bug-triggering function, which is identical to the function in the input program. The biggest difference between the existing approaches and RR-Reduce is most evident in its treatment of the functions that interact with the target function to create the necessary state to trigger the bug. In the existing approaches, every code that has been executed to create such a state is preserved. However, RR-Reduce just keeps the effects of the other code to make the jump happen, which yields a program with 48,862 bytes in terms of the target size. Hybrid-Reduce reduces the program even further to 33,358 bytes.
@@
Commanderkeen, which is 3.9MB in size, triggers a subtle wrong-code bug222Fixed by https://github.com/titzer/wizard-engine/commit/bc135ad in the JIT compiler. Wasm-reduce struggles the most, only reducing the input program to 1.3MB, which is hard for a human to manually debug. The bug is a subtle register allocation problem that occurs within the engine’s implementation of the call_indirect instruction. RR-Reduce selects the function where the behavior diverges as a target function and replaces the bodies of the functions that interact with the target function with replay functions, while deleting all the other functions, resulting in a reduced program of 782 bytes in terms of the target size. Hybrid-Reduce reduces the further down to 158 bytes, shown in Figure 1.
@@
Ffmpeg, which is 5.3MB in size, triggers a compiler-crash bug333Fixed by https://github.com/titzer/wizard-engine/commit/4e3e221. As this is a compiler-crash bug, most of the functions can be safely deleted except the bug-triggering function. Thus, wasm-reduce is effective in this case, reducing the input program to 479 bytes in 982 seconds. Hybrid-Reduce achieves an even smaller result of 45 bytes in only 150 seconds.
@@
Parquet, which is 1.7MB in size, triggers a trivial compiler-crash bug444Fixed by https://github.com/titzer/wizard-engine/commit/33ec201 in the Wizard engine. For this bug to be triggered, it is sufficient to contain a single memory declaration with 64-bit addresses. All other code can be removed as long as the Wasm module retains the offending memory declaration. Thus, wasm-reduce is effective, producing a reduced program that contains a single function with a single nop instruction in 216 seconds. Hybrid-Reduce achieves the same reduction but quicker; it takes only 25 seconds to obtain the reduced output.
@@ V-B Limitations
RR-Reduce uses simple string search over the engine’s output to heuristically identify a target function that triggers the bug. This heuristic fails for wrong-code bugs that do not manifest as observable changes in behavior. In such cases, RR-Reduce can fall back to the Dynamic set and All set at the cost of longer reduction time, yet still produce the same final minimized output. Another limitation is that RR-Reduce currently supports reduction of Wasm programs up to the Wasm 2.0 specification, excluding SIMD, due to the limitations of Wasm-R3. It also relies on the correctness and performance of Wasm-R3. Addressing more Wasm extensions and improving performance will require further engineering of the Wasm-R3 toolchain.
@@ VI Related Work
There are several dynamic analysis techniques for Wasm. Wasabi [19] designs and implements bytecode-level dynamic instrumentation for Wasm to enable diverse dynamic analyses. Wizard [31] supports non-intrusive instrumentation for Wasm by engine-level dynamic instrumentation. Wasm-R3 [12] is a record and replay technique for Wasm that enables the generation of executable, standalone Wasm benchmarks. Our work contributes to the field by utilizing the execution behavior of Wasm programs to improve Wasm program reduction.
@@ VII Conclusion
The artifact (source code, evaluation inputs/outputs, and documentation) is available at [18]

### v2

@@ Abstract
WebAssembly (Wasm) programs may trigger bugs in their engine implementations. To aid debugging, program reduction techniques try to produce a smaller variant of the input program that still triggers the bug. However, existing execution-unaware program reduction techniques struggle with large and complex Wasm programs, because they rely on static information and apply syntactic transformations, while ignoring the valuable information offered by the input program’s execution behavior. We present RR-Reduce and Hybrid-Reduce, novel execution-aware program reduction techniques that leverage execution behaviors via record and replay. RR-Reduce identifies a bug-triggering function as the target function, isolates that function from the rest of the program, and generates a reduced program that replays only the interactions between the target function and the rest of the program. Hybrid-Reduce combines a complementary execution-unaware reduction technique with RR-Reduce to further reduce program size. We evaluate RR-Reduce and Hybrid-Reduce on 28 Wasm programs that trigger a diverse set of bugs in three engines. On average, RR-Reduce reduces the programs to 1.20% of their original size in 14.5 minutes, which outperforms the state of the art by 33.15 in terms of reduction time. Hybrid-Reduce reduces the programs to 0.13% of their original size in 3.5 hours, which outperforms the state of the art by 3.42 in terms of reduced program size and 2.26 in terms of reduction time. We envision RR-Reduce as the go-to tool for rapid, on-demand debugging in minutes, and Hybrid-Reduce for scenarios where developers require the smallest possible programs.
@@ I Introduction
We release our tools and data as open source, for others to reproduce and build on our results: https://github.com/sola-st/rr-reduce.
@@ III-A Problem Statement
To the best of our knowledge, we are the first to formulate and address the execution-aware program reduction problem in this way. The key difference between the execution-aware program reduction problem and the standard program reduction problem is that the former considers an execution of the input program, while the latter does not.
@@ III-B Overview
Figure 3 shows an overview of RR-Reduce. Given an executable program as input, RR-Reduce first heuristically identifies functions in the program that may be critical for triggering the bug, which the approach considers as candidate target functions (Section III-D). Then, RR-Reduce iterates over the candidate target functions one by one, selecting each as the target function. To start an iteration, the approach splits the input program into two parts: one containing the target function and another containing all remaining functions (Section III-E). The resulting partitioned program is then passed to an existing record and replay technique, which records the execution and produces replay functions (Section III-F). Importantly, we configure the record and replay technique to focus the replay on replaying only the target function. That is, the remaining functions are either removed or replaced with replay functions that mimic those parts of the behavior of the original functions that are necessary to accurately replay the target function. Given the replay functions, RR-Reduce combines the target function with the replay functions to produce a candidate for the reduced program (Section III-G). Finally, the oracle, also given as input, checks whether the candidate program still triggers the bug (Section III-H). If so, and if the candidate program is smaller than the input program, RR-Reduce returns the candidate program as the reduced program. Otherwise, the approach repeats the iteration with a different target function.
@@ III-C Running Example
Figure 4 shows a running example of RR-Reduce for a single iteration. On the left, we have an input program that consists of three functions: a, b, and c. Suppose an execution of this input program triggers a wrong-code bug in a Wasm engine, which results in a stack trace that contains the c function. Because c is in the stack trace, RR-Reduce heuristically identifies c as the target function. Next, RR-Reduce splits the input program into two parts: one containing the target function c and another containing the remaining functions a and b. Then, RR-Reduce uses an existing record and replay technique, Wasm-R3 (Section II-B), to produce replay functions that accurately replay the execution of the target function c. This is achieved in three phases: record, reduce, and replay. The resulting replay functions mimic the behavior of the remaining functions a and b in the input program. Our approach combines these replay functions with the target function c to produce a candidate program, and checks using the oracle whether the candidate program still triggers the bug in the Wasm engine and is smaller than the input program.
@@ IV-A Experimental Setup
We measure the size of a Wasm program in terms of its code size, which is the size of the code section of the Wasm program. This excludes the size of the data section, which contains initialization data for the memory, and the size of the custom section, which contains debug symbols. Focusing on the code size is motivated by the fact that the code section is what engine developers typically focus on when debugging engine bugs. For RR-Reduce, we report two kinds of code sizes: the “All” code size, which includes all functions in the binary, and the “Target” code size, which is the size of the target function. The “Target” is more relevant for engine developers, as they only need to inspect the target function which is responsible for triggering the bug.
@@
We implement RR-Reduce through a combination of Python, Rust, and JavaScript, building on several libraries and tools in the Wasm ecosystem. The first step is to identify candidates for the target function (Section III-D). We use the existing tool, wasm-tools objdump [27], to compute the All set, a dynamic analysis based on the Wizard engine [13] to compute the Dynamic set, and implement the computation of the Heuristic set in Python. Next, our implementation splits the input program into two Wasm binaries (Section III-E), which we implement in Rust. In addition to two newly created Wasm binaries, our implementation also creates JavaScript glue code to dynamically link the two Wasm binaries. To resolve the circular dependency between the two binaries, the JavaScript code creates a closure that calls the target function, and instantiates the replay binary with this closure as an import. The record and replay step uses the existing Wasm-R3 [12] (commit hash 79b310b). Finally, our implementation combines the target function with the replay functions (Section III-G) using the wasm-merge tool [28]. To speed up the reduction process, our implementation tries to reduce the input program for different target functions in parallel. We ran all experiments on an Ubuntu 24.04 LTS system with an Intel Core i9-13900K (32 logical cores) and 192 GB of DRAM.
@@
We evaluate the effectiveness of our approach by applying RR-Reduce and Hybrid-Reduce, as well as the two baselines, to the Wasm programs in Table I. We run each technique either until it terminates or a timeout of 24 hours is reached. We run eight reduction tasks in parallel, allowing each task to use a maximum of four logical cores. This ensures a fair comparison between wasm-reduce, RR-Reduce, and Hybrid-Reduce, which all exploit parallelism. 111Wasm-shrink does not exploit parallelism.
@@
The real strength of RR-Reduce is revealed when it is combined with existing program reduction techniques into a hybrid approach. Hybrid-Reduce reduces the input program to 0.13% of its original size, which is a 3.42 improvement over the current state of the art, wasm-reduce. Hybrid-Reduce achieves this by giving the best result in 23 cases, of which six are ties with the current state of the art. The benefit of the Hybrid-Reduce is not only in its better average performance, but also in its effectiveness on programs where the current state of the art struggles. Table III highlights the difference. Specifically, wasm-shrink produces programs larger than 100KB in 15 cases, and wasm-reduce does so in three cases. In contrast, all Hybrid-Reduce-reduced programs are well below 100KB, and often even further.
@@
IV-C RQ2. Efficiency
RR-Reduce and Hybrid-Reduce are able to achieve this speedup due to their heuristic identification of the target function. Among the 28 programs in the evaluation set, the target function is identified in the Heuristic set in 13 cases, the Dynamic set in 6 cases, and the All set in 9 cases. As the average size of the Heuristic set is 2.68, the Dynamic set is 277.04, and the All set is 1,995.57, identifying the target function in the Heuristic set greatly reduces the search space. Even if the heuristic fails, RR-Reduce can fall back to the Dynamic or All sets, losing some efficiency but without affecting its effectiveness.
IV-D RQ3. Qualitative Analysis
@@
Boa, which is 5.2MB in size, triggers a wrong-code bug222Fixed by https://github.com/titzer/wizard-engine/commit/6d2b057 that gets triggered when the interpreter tries to interpret a jump with an offset bigger than 32,768 bytes. Wasm-reduce reduces the input program to 317KB, which is still too large to manually debug. It also contains 87 functions besides the function where the jump happens, which further complicates debugging. Wasm-reduce cannot simply remove these functions because other functions are necessary for the jump to happen. In contrast, RR-Reduce takes a different approach: It deletes the bodies of 453 executed functions and replaces the bodies of eight functions with replay functions. Among the nine functions that are executed in the reduced program, seven are short functions that update the state, and one is the entry to the reduced program, which calls the bug-triggering functions. The remaining one is the bug-triggering function, which is identical to the function in the input program. The biggest difference between the existing approaches and RR-Reduce is most evident in its treatment of the functions that interact with the target function to create the necessary state to trigger the bug. In the existing approaches, every code that has been executed to create such a state is preserved. However, RR-Reduce just keeps the effects of the other code to make the jump happen, which yields a program with 48,862 bytes in terms of the target size. Hybrid-Reduce reduces the program even further to 33,358 bytes.
@@
Commanderkeen, which is 3.9MB in size, triggers a subtle wrong-code bug333Fixed by https://github.com/titzer/wizard-engine/commit/bc135ad in the JIT compiler. Wasm-reduce struggles the most, only reducing the input program to 1.3MB, which is hard for a human to manually debug. The bug is a subtle register allocation problem that occurs within the engine’s implementation of the call_indirect instruction. RR-Reduce selects the function where the behavior diverges as a target function and replaces the bodies of the functions that interact with the target function with replay functions, while deleting all the other functions, resulting in a reduced program of 782 bytes in terms of the target size. Hybrid-Reduce reduces the further down to 158 bytes, shown in Figure 1.
@@
Ffmpeg, which is 5.3MB in size, triggers a compiler-crash bug444Fixed by https://github.com/titzer/wizard-engine/commit/4e3e221. As this is a compiler-crash bug, most of the functions can be safely deleted except the bug-triggering function. Thus, wasm-reduce is effective in this case, reducing the input program to 479 bytes in 982 seconds. Hybrid-Reduce achieves an even smaller result of 45 bytes in only 150 seconds.
@@
Parquet, which is 1.7MB in size, triggers a trivial compiler-crash bug555Fixed by https://github.com/titzer/wizard-engine/commit/33ec201 in the Wizard engine. For this bug to be triggered, it is sufficient to contain a single memory declaration with 64-bit addresses. All other code can be removed as long as the Wasm module retains the offending memory declaration. Thus, wasm-reduce is effective, producing a reduced program that contains a single function with a single nop instruction in 216 seconds. Hybrid-Reduce achieves the same reduction but quicker; it takes only 25 seconds to obtain the reduced output.
@@ V-B Limitations
RR-Reduce uses simple string search over the engine’s output to heuristically identify a target function that triggers the bug. This heuristic can be effective for engine crash bugs and wrong-code bugs that lead to a runtime error. It is not effective for wrong-code bugs that do not manifest as observable changes in behavior. In such cases, RR-Reduce can fall back to the Dynamic set and All set at the cost of longer reduction time, yet still produce the same final minimized output. Another limitation is that RR-Reduce currently supports reduction of Wasm programs up to the Wasm 2.0 specification, excluding SIMD, due to the limitations of Wasm-R3. It also relies on the correctness and performance of Wasm-R3. Addressing more Wasm extensions and improving performance will require further engineering of the Wasm-R3 toolchain.
V-C Generalization to Other Programming Languages
Conceptually, RR-Reduce can be applied to other programming languages. There are a few conditions for this to be possible: (1) A record and replay implementation must exist. (2) The record and replay system must support selective record and replay. (3) The output of selective record and replay must be a regular program.
@@ VI Related Work
There are several dynamic analysis techniques for Wasm. Wasabi [18] designs and implements bytecode-level dynamic instrumentation for Wasm to enable diverse dynamic analyses. Wizard [30] supports non-intrusive instrumentation for Wasm by engine-level dynamic instrumentation. Wasm-R3 [12] is a record and replay technique for Wasm that enables the generation of executable, standalone Wasm benchmarks. Wemby [31] utilizes dynamic analysis to detect memory corruption bugs in Wasm. Our work contributes to the field by utilizing the execution behavior of Wasm programs to improve Wasm program reduction.
@@ VII Conclusion
The artifact (source code, benchmark, evaluation outputs, and documentation) is available at https://github.com/sola-st/rr-reduce.
Acknowledgements
We thank Laurence Tratt and anonymous reviewers for their helpful feedback. This work was partially supported by NSF award #2148301, the WebAssembly Research Center, the National Research Foundation of Korea (NRF) (2022R1A2C2003660 and 2021R1A5A1021944), an Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (2024-00337703), Samsung Electronics Co., Ltd., by the European Research Council (ERC, grant agreements 851895 and 101155832), and by the German Research Foundation within the DeMoCo project.